Suppose two lumps of clay approach each other from opposite



directions in deep space. Lump number 1 masses 4.00 kg and travels at



5.00 m/s. Lump number 2 masses 10.0 kg. When the lumps collide, no



material flies off. The lumps stick together , and the resulting object is



stationary relative to you, the observer. How fast was lump number 2



moving before the collision relative to you, the observer?

1
Expert's answer
2021-12-02T10:08:54-0500

Let us calculate the momenta of lumps. Before the collision the first lump had a momentum of "m_1v_1 = 4.00\,\mathrm{kg}\cdot5.00\mathrm{m/s}=20\,\mathrm{kg\cdot m/s}."


The second momentum is "m_2v_2 = 10.0\,\mathrm{kg}\cdot v_2"

and the direction of this momentum is opposite to the direction of the first momentum.


Due to the law of conservation of momenta we have

"m_1v_1 - m_2v_2 =(m_1+m_2)v,"

where we take the opposite directions of motion into account and where "v=0".

Therefore,

"m_1v_1=m_2v_2, v_2 = 2\,\text{m/s}."